OpenStudy (lgbasallote):

how to integrate \[\int \frac{2+v}{1+v^2}\]

OpenStudy (blockcolder):

Separate into 2 fractions: \[\frac{2}{1+v^2}+\frac{v}{1+v^2}\] then integrate separately.

OpenStudy (anonymous):

seperate it into 2 integ question first is intg (2/(1+v2)) and the other is 1/2 intg(2v/(1+v2)) and thus the ans is 2tan-1 v +1/2 ln|1+v2| + c

OpenStudy (lgbasallote):

@blockcolder how do you integrate 2/(1+v)^2

OpenStudy (blockcolder):

Don't you mean 2/(1+v^2)?

OpenStudy (lgbasallote):

yeah

OpenStudy (blockcolder):

Here's a rule for safe-keeping: \[\int \frac{1}{1+x^2} dx=\tan^{-1}x+C\]

OpenStudy (lgbasallote):

oh..yeah of course *facepalm*